OpenStudy (anonymous):

Part 1: Solve the inequality -3x + 4 x + 8 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1.

OpenStudy (anonymous):

@ritez

OpenStudy (anonymous):

-3x + 4 <= x + 8 subtract x both the sides -4x + 4 <= 8 now subtract 8 both the sides -4x -4 <= 0 -4( x + 1) <= 0 now multiplying with negative sign both the sides 4 ( x + 1) >= 0 x + 1 >= 0 x >= -1